1. If a 25.0 L sample of gas at 0C is heated to 35 0C, what is the new sample of the gas at constant pressure?
1
Expert's answer
2021-01-22T06:26:38-0500

AccordingtogaslawV1/T1=V2/T2V2=V1∗T2/T1=25∗298/273=27.3l